Sophie was 28 years old when her daughter Nichole was born. In how many years will Sophie be two years less than six times Nicholes age?
Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13200)



Let x be Nichole's age when Sophie is 2 less than 6 times Nichole's age.

Then of course Sophie's age then will be 6x-2; the DIFFERENCE in their ages will be (6x-2)-x = 5x-2.

The difference in their ages will always be 28 years. So

5x-2+=+28

Finish the problem using basic algebra.

Answer by ikleyn(52786)

.

From the condition,  Sophie is 28 years older than her daughter Nichole.


The question asks: How old  Nichole will be when Sophie is two years less than 6 times Nichole's age ?


So, let "x" be the Nichole age at this time.

Then Sophie age is  (x+28) years.


At the questioned time, this equation will be held


    x + 28 = 6x - 2.


Simplify and solve for x


   28 + 2 = 6x - x

   30     = 5x

    x     = 30/5 = 6.


ANSWER.  In 6 years.


CHECK.   Sophie will be 6 years old;  Nichole will be 6+28 = 34, which is 6*6-2 = 34 years.    ! Correct !

Solved.